# Smooth Tests for Normality in ANOVA

### v3

@@ Appendix A Appendix
where the second last line is followed by (A.22), and $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=c_{1k}$ under $H_{0}$. The last line is followed by $N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=o_{p}(1)$ using LLN. For $D_{N2}$, under $H_{0}$, we have
$D_{N2}=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}\left\{\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\hat{\mu}-\mu\right)^{2}\right$.
$\left.-2\left(\hat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{2}}{N_{j}\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}\left(e_{i_{2}j}^{2}-1\right)\right\}$
$+\left(\hat{\mu}-\mu\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}$
$-2\left(\hat{\mu}-\mu\right)\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}}{N_{j}\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}\right\}$
@@
$+\left(\hat{\mu}-\mu\right)^{2}\frac{1}{2N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}\left(1+o_{p}(1)\right)$
$-\left(\hat{\mu}-\mu\right)\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}\sigma_{j}}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}\right\}$
@@
following (A.20), (A.21), (A.23), (A.24) and (A.25), as well as $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=c_{2k}$ under $H_{0}$.
@@
$|R_{N1}|\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}_{j}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\hat{\sigma}_{j}\sigma_{j}(\hat{\sigma}_{j}+\sigma_{j})}\right)^{2}\equiv R_{N1}^{A}+R_{N1}^{B}.$
@@
$R_{N1}^{A}=\frac{(\hat{\mu}-\mu)^{2}}{N}\sum_{j=1}^{J}\frac{N_{j}}{\hat{\sigma}_{j}^{2}}\asymp\frac{1}{N}O_{p}\left(\frac{1}{N}\right)O_{p}(N)=O_{p}\left(\frac{1}{N}\right)=o_{p}\left(\frac{1}{\sqrt{N}}\right),$
@@
$\sum_{j=1}^{J}\frac{N_{j}}{\hat{\sigma}_{j}^{2}}=\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}^{2}}(1+o_{p}(1))=O_{p}(N).$
@@
$R_{N1}^{B}\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\left(\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\hat{\mu}-\mu\right)^{2}-2\left(\hat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\left(\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)+\left(\hat{\mu}-\mu\right)^{4}\left(\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
$+4\left(\hat{\mu}-\mu\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{4}}{N_{j}^{2}}\left(\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
$\equiv$ $R_{N1}^{B,1}+\left(\hat{\mu}-\mu\right)^{4}R_{N1}^{B,2}+4\left(\hat{\mu}-\mu\right)^{2}R_{N1}^{B,3}.$
@@
Note that $\mathrm{E}[(e_{i_{1}j}^{2}-1)(e_{i_{2}j}^{2}-1)e_{i_{3}j}^{2}]=\mathrm{E}[(e^{2}-1)^{2}e^{2}]\mathbbm{1}(i_{1}=i_{2}=i_{3})+\mathrm{E}[(e^{2}-1)^{2}]\mathbbm{1}(i_{1}=i_{2}\neq i_{3})$. Then,
$\mathrm{E}\left[R_{N1}^{B,1}\right]$ $=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\left(N_{j}\mathrm{E}\left[(e^{2}-1)^{2}e^{2}\right]+N_{j}(N_{j}-1)\mathrm{E}\left[(e^{2}-1)^{2}\right]\right)$
@@
which implies $R_{N1}^{B,1}=o_{p}(N^{-1/2})$. Similarly, for $R_{N1}^{B,3}$, by $\mathrm{E}[e_{i_{1}j}e_{i_{2}j}e_{i_{3}j}^{2}]=\mathrm{E}[e^{4}]\mathbbm{1}(i_{1}=i_{2})+\mathbbm{1}(i_{1}=i_{2}\neq i_{3})$, we have
$\mathrm{E}\left[R_{N1}^{B,3}\right]=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{4}}{N_{j}^{2}}\left(N_{j}\mathrm{E}\left[e^{4}\right]+N_{j}(N_{j}-1)\right)=O\left(\frac{J}{N}\right),$
@@
$\mathrm{E}\left[R_{N1}^{B,2}\right]=\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\sum_{i=1}^{N_{j}}e_{ij}^{2}\right]=\frac{1}{N}\sum_{j=1}^{J}N_{j}\sigma_{j}^{2}=O(1),$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)^{4}$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}_{j}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)}{\hat{\sigma}_{j}\sigma_{j}(\hat{\sigma}_{j}+\sigma_{j})}\right)^{4}$
$\asymp\left(\hat{\mu}-\mu\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{4}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}$
@@
$\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{4}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}$
$\leq$ $\left\{\max_{1\leq j\leq J}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\right\}\left\{\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}\right\}$
$=$ $o_{p}\left(\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}\right)$
@@
where we use $\hat{\sigma}^{2}_{j}-\sigma^{2}_{j}=o_{p}(1)$ from (A.21), and the last step can be obtained through similar arguments above. Besides, $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\tilde{\varepsilon}_{ij}^{2}/\tilde{\sigma}_{j}^{2}=O_{p}(1)$ can be derived from the procedure in (A.9) analogously. Then we have
$\left|R_{N2}\right|\lesssim\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}_{j}^{2}}\right]^{1/2}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)^{4}\right]^{1/2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$

### v4

@@ A.5 Proof of Theorem 7
where the second last line is followed by (A.22), and $\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=c_{1k}$ under $H_{0}$. The last line is followed by $N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=o_{p}(1)$ using LLN. For $D_{N2}$, under $H_{0}$, we have
$D_{N2}=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\widehat{\sigma}_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\widehat{\sigma}_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}\left\{\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\widehat{\mu}-\mu\right)^{2}\right$.
$\left.-2\left(\widehat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{2}}{N_{j}\widehat{\sigma}_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}\left(e_{i_{2}j}^{2}-1\right)\right\}$
$+\left(\widehat{\mu}-\mu\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\widehat{\sigma}_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}$
$-2\left(\widehat{\mu}-\mu\right)\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}}{N_{j}\widehat{\sigma}_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}\right\}$
@@
$+\left(\widehat{\mu}-\mu\right)^{2}\frac{1}{2N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}\left(1+o_{p}(1)\right)$
$-\left(\widehat{\mu}-\mu\right)\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}\sigma_{j}}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}\right\}$
@@
following (A.20), (A.21), (A.23), (A.24) and (A.25), as well as $\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=c_{2k}$ under $H_{0}$.
@@
$|R_{N1}|\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}_{j}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\widehat{\sigma}_{j}\sigma_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\right)^{2}\equiv R_{N1}^{A}+R_{N1}^{B}.$
@@
$R_{N1}^{A}=\frac{(\widehat{\mu}-\mu)^{2}}{N}\sum_{j=1}^{J}\frac{N_{j}}{\widehat{\sigma}_{j}^{2}}\asymp\frac{1}{N}O_{p}\left(\frac{1}{N}\right)O_{p}(N)=O_{p}\left(\frac{1}{N}\right)=o_{p}\left(\frac{1}{\sqrt{N}}\right),$
@@
$\sum_{j=1}^{J}\frac{N_{j}}{\widehat{\sigma}_{j}^{2}}=\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}^{2}}(1+o_{p}(1))=O_{p}(N).$
@@
$R_{N1}^{B}\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\left(\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\widehat{\mu}-\mu\right)^{2}-2\left(\widehat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\left(\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)+\left(\widehat{\mu}-\mu\right)^{4}\left(\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
$+4\left(\widehat{\mu}-\mu\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{4}}{N_{j}^{2}}\left(\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
$\equiv$ $R_{N1}^{B,1}+\left(\widehat{\mu}-\mu\right)^{4}R_{N1}^{B,2}+4\left(\widehat{\mu}-\mu\right)^{2}R_{N1}^{B,3}.$
@@
Note that $\mathbb{E}[(e_{i_{1}j}^{2}-1)(e_{i_{2}j}^{2}-1)e_{i_{3}j}^{2}]=\mathbb{E}[(e^{2}-1)^{2}e^{2}]\mathbbm{1}(i_{1}=i_{2}=i_{3})+\mathbb{E}[(e^{2}-1)^{2}]\mathbbm{1}(i_{1}=i_{2}\neq i_{3})$. Then,
$\mathbb{E}\left[R_{N1}^{B,1}\right]$ $=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\left(N_{j}\mathbb{E}\left[(e^{2}-1)^{2}e^{2}\right]+N_{j}(N_{j}-1)\mathbb{E}\left[(e^{2}-1)^{2}\right]\right)$
@@
which implies $R_{N1}^{B,1}=o_{p}(N^{-1/2})$. Similarly, for $R_{N1}^{B,3}$, by $\mathbb{E}[e_{i_{1}j}e_{i_{2}j}e_{i_{3}j}^{2}]=\mathbb{E}[e^{4}]\mathbbm{1}(i_{1}=i_{2})+\mathbbm{1}(i_{1}=i_{2}\neq i_{3})$, we have
$\mathbb{E}\left[R_{N1}^{B,3}\right]=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{4}}{N_{j}^{2}}\left(N_{j}\mathbb{E}\left[e^{4}\right]+N_{j}(N_{j}-1)\right)=O\left(\frac{J}{N}\right),$
@@
$\mathbb{E}\left[R_{N1}^{B,2}\right]=\mathbb{E}\left[\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\sum_{i=1}^{N_{j}}e_{ij}^{2}\right]=\frac{1}{N}\sum_{j=1}^{J}N_{j}\sigma_{j}^{2}=O(1),$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)^{4}$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}_{j}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)}{\widehat{\sigma}_{j}\sigma_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\right)^{4}$
$\asymp\left(\widehat{\mu}-\mu\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{4}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}$
@@
$\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{4}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}$
$\leq$ $\left\{\max_{1\leq j\leq J}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\right\}\left\{\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}\right\}$
$=$ $o_{p}\left(\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}\right)$
@@
where we use $\widehat{\sigma}^{2}_{j}-\sigma^{2}_{j}=o_{p}(1)$ from (A.21), and the last step can be obtained through similar arguments above. Besides, $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\widetilde{\varepsilon}_{ij}^{2}/\widetilde{\sigma}_{j}^{2}=O_{p}(1)$ can be derived from the procedure in (A.9) analogously. Then we have
$\left|R_{N2}\right|\lesssim\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\widetilde{\varepsilon}_{ij}^{2}}{\widetilde{\sigma}_{j}^{2}}\right]^{1/2}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)^{4}\right]^{1/2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$